BAT Industries is a manufacturer of orange juice in Jamaica. In 2022 it sold a bottle of orange juice to wholesalers for $100 and its variable costs is $40 and fixed costs is $100,000 monthly:

 

 

Using the information provided. 

 

i. How many bottles of orange juice must the company sell in order to break even and what is the breakeven revenue? 

 

 

 

ii. If the company sold 1,500,000 bottles of orange juice for the year 2022, what was its profit? 

 

 

 

iii. Due to inflation, in 2023 fixed costs are expected to rise to $175,000 monthly. What will be the NEW break-even quantity and revenue? 

 

 

 

iiii. If the company will sell the same number of orange juice as it did in 2022 and it wants to make the same profit as in 20222, what price should they sell each bottle of orange Juice for in 2023? 

 

 

 



1
Expert's answer
2022-03-15T17:14:05-0400

1.) Find contribution margin=price - Variables=100-40=60

"Bottlers=\frac{100,000}{60}=1,667"

2.) Profit=Revenue- Cost=1,500,000(100-40)-100000=$89,900,000

3.) New break even

"Bottler=\frac{175000}{60}=2917"

4.) Produce between 1000 and 3000